OpenStudy (anonymous):

HELP PLEASE!!! In Gloria's nursery, 5/7 of the plants were orange trees. An orchard owner bought 2/5 of all the orange trees from the nursery. Show the steps that can be used to calculate the fraction of the total number of plants in the nursery that the orchard owner bought.

OpenStudy (anonymous):

5/7 = 25/35 2/5 = 14/35